# Learning from Imperfect Human Feedback: a Tale from Corruption-Robust $\text{Dueling}^{*}$

### v1

@@ Abstract
This paper studies Learning from Imperfect Human Feedback (LIHF), motivated by humans’ potential irrationality or imperfect perception of true preference. We revisit the classic dueling bandit problem as a model of learning from comparative human feedback, and enrich it by casting the imperfection in human feedback as agnostic corruption to user utilities. We start by identifying the fundamental limits of LIHF and prove a regret lower bound of $\Omega(\max\{T^{1/2},C\})$, even when the total corruption $C$ is known and when the corruption decays gracefully over time (i.e., user feedback becomes increasingly more accurate). We then turn to design robust algorithms applicable in real-world scenarios with arbitrary corruption and unknown $C$. Our key finding is that gradient-based algorithms enjoy a smooth efficiency-robustness tradeoff under corruption by varying their learning rates. Specifically, under general concave user utility, Dueling Bandit Gradient Descent (DBGD) of Yue and Joachims (2009) can be tuned to achieve regret $O(T^{1-\alpha}+T^{\alpha}C)$ for any given parameter $\alpha\in(0,\frac{1}{4}]$. Additionally, this result enables us to pin down the regret lower bound of the standard DBGD (the $\alpha=1/4$ case) as $\Omega(T^{3/4})$ for the first time, to the best of our knowledge. For strongly concave user utility we show a better tradeoff: there is an algorithm that achieves $O(T^{\alpha}+T^{\frac{1}{2}(1-\alpha)}C)$ for any given $\alpha\in[\frac{1}{2},1)$. Our theoretical insights are corroborated by extensive experiments on real-world recommendation data.
@@ 1 Introduction
Many real-world problems, such as personalized recommendation (Yue and Joachims, 2009; Immorlica et al., 2020; Yao et al., 2022a) and fine-tuning generative models (Bai et al., 2022; Casper et al., 2023; Han et al., 2024), require learning from human feedback. Expressing preferences as numerical values or concrete functions is generally challenging for humans. Therefore, an approach that has achieved significant real-world success is to learn humans’ preferences by eliciting their comparative feedback between two options (Bai et al., 2022). A natural theoretical framework capturing such learning task is the seminal dueling bandits framework introduced by Yue and Joachims (2009). The dueling bandit problem features an sequential online learning problem, during which a pair of actions are selected at each round and only their comparison result is revealed to the learner. The comparison outcome between two actions is modeled using a utility function of actions, alongside a link function that determines the probability of each action winning based on their utility difference. This modeling approach is termed as utility-based dueling bandit and achieves great success in generating summaries closely aligned with human preferences (Stiennon et al., 2020).
@@
However, human feedback is often imperfect, a crucial factor that has been largely overlooked in previous studies of dueling bandit. Ample behavioral studies show that humans often refine their preferences through interactions with the system, navigating a complex tradeoff between exploration and exploitation (Wilson et al., 2014; Cohen et al., 2007). Furthermore, their response errors may exhibit systematic patterns rather than merely random noise. As we are all aware of, humans are not always rational (Posner, 1997) neither perfectly know our preferences (Pu et al., 2012).
In this paper, we study learning from imperfect human feedback (LIHF) by casting the imperfection in the feedback as certain form of corruption to the utilities. This leads us to study a natural variant of the continuous dueling bandit with corrupted utilities. Notably, our study differs from recent robust algorithms designed for corrupted dueling bandits (Di et al., 2024; Saha and Gaillard, 2022; Komiyama et al., 2015) in three key aspects. Firstly, these previous algorithms all focus on a finite action space whereas, to the best of our knowledge, ours is the first to investigate a robust dueling bandit algorithm under corruption for continuous action space. This focus is motivated by the fact that many real-world applications have continuous or extremely dense action spaces, such as token embeddings or content recommendations (Chen et al., 2021, 2019). Second, our algorithms are designed for agnostic corruption without knowing the corruption levels, whereas Di et al. (2024) requires knowledge of the total corruption amount. This focus on agnostic corruption is motivated by realistic considerations, because systems often cannot accurately discern the irrational or imperfect nature of human feedback. Third, our corruption models are very different from the model studied by Di et al. (2024); Saha and Gaillard (2022); Komiyama et al. (2015). The corruption in their model contaminates feedback by flipping the realized outcomes of the duels, whereas the corruption in our model is on utilities which consequently only alter the winning probabilities of the duels. This difference is due to our different motivation of corruption, which is to capture imperfect human behavior while not intentional adversarial attack (see more illustration in Section 3).
Our Contributions. By modeling LIHF as a continuous dueling bandit under agnostic corruption, we start by characterizing its intrinsic hardness. We prove a regret lower bound of $\Omega(d\max\{\sqrt{T},C\})$, even when the total corruption $C$ is known and when corruption decays gradually over time. Additionally, we prove that this lower bound result is tight by exhibiting an algorithm that indeed achieves a matching upper bound for the aforementioned setting. Next, we turn to design robust algorithms that are applicable to arbitrary and agnostic corruption with unknown $C$. Our key finding is that gradient-based algorithms offer a smooth efficiency-robustness tradeoff under corruption by carefully tuning their learning rates. Specifically, let $d$ denote the problem dimension. For any given $\alpha\in(0,1/4]$, Dueling Bandit Gradient Descent (DBGD) (Yue and Joachims, 2009) with carefully chosen learning rate (that depends on $\alpha$) can achieve a regret of $O(\sqrt{d}T^{1-\alpha}+\sqrt{d}T^{\alpha}C)$ under arbitrary concave user utility. That is, by decreasing $\alpha$ from $1/4$ to 00, the regret under no corruption (i.e., learning efficiency) decays from $\sqrt{d}T^{3/4}$ to $\sqrt{d}T$ but the algorithm’s tolerance to total amount of corruption increases from $O(T^{3/4})$ to $O(T)$ (i.e., more robustness). Perhaps surprisingly, this result turns out to allow us to pin down a regret lower bound for the standard version of DBGD (the $\alpha=1/4$ case) as $\Omega(T^{3/4})$, marking the first tight lower bound result for the DBGD algorithm to the best of our knowledge. For strongly concave user utility, we show that an improved tradeoff is possible: Noisy Comparison based Stochastic Mirror Descent (NC-SMD) (Kumagai, 2017) can be refined to achieve a regret of $O(dT^{\alpha}+\sqrt{d}T^{\frac{1-\alpha}{2}}C+dC)$ for any given $\alpha\in[1/2,1)$. We remark that similar robustness-accuracy tradeoff has been studied classification problem with adversarial examples, both empirically (Tsipras et al., 2018) and theoretically (Zhang et al., 2019). However, to the best of our knowledge, this is the first time such tradeoff is formally analyzed for online learning. Finally, we conduct extensive experiments on real-world recommendation data to validate our theoretical insights.
2 Related Work
Hardness of Learning from Imperfect Human Feedback: To design algorithms robust to adversarial corruption, it is crucial to understand the intrinsic hardness of learning from corrupted feedback. Lattimore and Szepesvári (2020) demonstrates that the minimax regret lower bound for stochastic linear bandits is $\Omega(d\sqrt{T})$. Utilizing the regret lower bound for bandit convex optimization (Shamir, 2013), Kumagai (2017) illustrates that dueling bandits with a strongly concave utility function achieve a regret lower bound of $\Omega(d\sqrt{T})$. In scenarios of adversarial corruption, Bogunovic et al. (2020) proves that any algorithm faces regret at least $\Omega(dC)$ for stochastic linear bandits. However, currently there is no literature that has explored the intrinsic hardness of learning from corruption which decays gradually over time (Wilson et al., 2014; Kumagai, 2017; Cohen et al., 2007), despite its prevalence in learning from imperfect human feedback. It still remains unknown whether learning from such corruption is fundamentally easier than learning from arbitrary corruption.
Stochastic Bandits with Unknown Corruption: Applying the principle of optimism in face of uncertainty and constructing estimators involving uncertainty weighting, algorithms devised by He et al. (2022) and Ye et al. (2024a) attain nearly optimal regret bounds for both linear and nonlinear contextual bandits in presence of adversarial corruption. By leveraging the concept of phase elimination, Bogunovic et al. (2022) introduces a robust algorithm capable of handling corruption in rewards drawn from Gaussian processes. However, the designs of these algorithms require explicit knowledge of the total corruption level $C$ as an input. Gupta et al. (2019) employ novel sampling strategy to channel resources to arms with better performance and craft algorithms robust to agnostic corruption, yet restricted to finite action sets. In scenarios involving continuous action set with unknown corruption level, even though algorithms designed by Ding et al. (2022) and Bogunovic et al. (2020) have sublinear regret guarantee, the regret is at least $C$ times worse than the regret without corruption. It remains as an open problem of whether better regret guarantee is achievable—that is, affording unknown corruption level more than $\Omega(\sqrt{T})$ while maintaining $O(\sqrt{T})$ regret guarantee in scenario without corruption.
Continuous Dueling Bandits with Unknown Corruption: In addition to stochastic bandits, there exists an extensive body of literature studying the dueling bandit problem with unknown corruption, given its prevalence in real-world application (Bengs et al., 2021b; Sui et al., 2018). The literature on dueling bandits with adversarial corruption has been initially explored by Agarwal et al. (2021) and subsequently enhanced by Saha and Gaillard (2022), which achieves optimality in term of $C$. However, their work is solely applicable to finite action sets, which may not be feasible in many real-world scenarios that have continuous or extremely dense action spaces, such as token embeddings or content recommendations (Chen et al., 2021, 2019). Ailon et al. (2014) introduced a reduction method that translates existing results from conventional Multi-Armed Bandit algorithms to Dueling Bandits. This reduction method is applicable to both finite and continuous action spaces, albeit requiring a linear link function. DBGD (Yue and Joachims, 2009) and NC-SMD (Kumagai, 2017) enjoy flexibility in choosing the link function but only work for continuous dueling bandits. It still remains as an open problem for developing provably efficient algorithm capable of learning from corrupted dueling feedback with unknown corruption in continuous action space. This serves as the driving force behind our theoretical analysis.
3 The Problem of Learning from Imperfect Human Feedback (LIHF)
@@
Conventional Notation. Throughout the paper, for a positive integer $T$, we use $[T]$ to denote $\{1,2,\ldots,T\}$. We use $\mathcal{P}_{\mathcal{A}}(a)$ to denote the projection of action $a$ on $\mathcal{A}$. We use standard asymptotic notations including $O(\cdot)$, $\Omega(\cdot)$, $\Theta(\cdot)$. We use $\tilde{O}(\cdot)$, $\tilde{\Omega}(\cdot)$, $\tilde{\Theta}(\cdot)$ to hide logarithmic factors. We use $\|\cdot\|_{\infty}$ to define infinity norm. $\text{Reg}_{T}$ represents regret with dueling feedback over $T$ rounds.
To capture learning from imperfect (comparative) human feedback, we now introduce a natural variant of the seminal dueling bandit framework (Yue and Joachims, 2009) by integrating the “imperfections” in human feedback as certain corruption to human agents’ feedback. An immediate critique one might have is that arbitrarily adversarial corruption seems too pessimistic as a way to model humans’ inaccuracies which are often caused by uncertainty or ignorance, rather than targeted attacks. While this may appear true at the first glance, we provide strong theoretical and empirical evidences that corruption caused by uncertainties generally is as hard as arbitrary corruption. For such reasons, we coin the following problem Learning from Imperfect Human Feedback (LIHF).
@@
Basics of Dueling Bandits and Assumptions. We consider the standard dueling bandit framework with continuous action set $\mathcal{A}\subset\mathbb{R}^{d}$ (Yue and Joachims, 2009). At each round $t\in[T]$, the learner (henceforth denoted as $\mathcal{L}$) chooses two actions $a_{t},a^{\prime}_{t}$ to present to some human agent, henceforth denoted as the “user”. The user receives utility $\mu(a_{t}),\mu(a^{\prime}_{t})$ from the two actions respectively, and will pick one of these actions following a link function $\sigma(\cdot)$. That is, in absence of any corruption, the user picks action $a_{t}$ with probability $\sigma(\mu(a_{t})-\mu(a^{\prime}_{t}))$, and action $a^{\prime}_{t}$ otherwise. As clear from its definition, this choice probability depends only on the utility difference $\mu(a_{t})-\mu(a^{\prime}_{t})$. Notably, user’s randomized action does carry information about underlying utility function $\mu$, which is why the action is also called the “human feedback” (Bengs et al., 2021a; Maystre and Grossglauser, 2017; Ailon et al., 2014). Following standard assumptions in this literature (Kumagai, 2017; Yue and Joachims, 2009), we also assume
@@
    the continuous action space $\mathcal{A}$ contains the origin, is compact, convex and is contained in a $d$-dimensional ball of radius $R$;
@@
    the utility function $\mu:\mathcal{A}\rightarrow\mathbb{R}$ is concave, differentiable, $L_{\mu}$-Lipschitz and admits unique optimal within set $\mathcal{A}$, denoted as $a^{*}:=\arg\max_{a\in\mathcal{A}}\mu(a)$;
@@
    the link function $\sigma:\mathbb{R}\rightarrow[0,1]$ is rotational-symmetric, $L_{\sigma}$-Lipschitz and its first derivative $\sigma^{\prime}$ is $L_{2}$-Lipschitz.
@@
For convenience, we further let $L=L_{\sigma}L_{\mu}$. Notably, some previous works restrict link function to specifically be the logistic link function (Di et al., 2024; Xu et al., 2024; Saha, 2021). However, our results shall be applicable to any link functions that satisfy the natural assumptions above.
Inaccuracies in Human Feedback as Agnostic Corruption. Extending the standard dueling bandit framework above, we introduce a corruption term $c_{t}(a_{t},a^{\prime}_{t})$ to the utility difference in order to model potential inaccuracy in the user’s perception of underlying true utilities. For all our positive results, we always work with the situation where this corruption is agnostic in the sense that the learner does not know the value of each $c_{t}$ neither their accumulated sum. Such agnostic corruption is natural because in reality it is very difficult to truly estimate how imperfect a user is, hence we aim to develop algorithms whose execution is agnostic to such parameters. Let $a\succ a^{\prime}$ denote the event that the user chooses $a$ over $a^{\prime}$. Hence in our LIHF problem, the probability of event $a\succ a^{\prime}$ in presence of corruption $c(a,a^{\prime})$, denoted as ^$\hat{\mathbb{P}}(a\succ a^{\prime})$, can be expressed as follows ^$\hat{\mathbb{P}}(a\succ a^{\prime})=\sigma(\mu(a)-\mu(a^{\prime})+c(a,a^{\prime}))$. We further let ^$\hat{\mathcal{F}}(a_{t},a^{\prime}_{t})$ denote user’s preferential feedback under corruption, coined corrupted dueling feedback. Mathematically, ^$\hat{\mathcal{F}}(a_{t},a^{\prime}_{t})$ follows a binomial distribution with mean ^$\hat{\mathbb{P}}(a_{t}\succ a^{\prime}_{t})$, expressed by the following equation.
@@
^$\hat{\mathcal{F}}(a_{t},a^{\prime}_{t}):=\begin{cases}1\hskip 2.84526pt\text{w.p.}\hskip 2.84526pt\hat{\mathbb{P}}(a_{t}\succ a^{\prime}_{t})\\ 0\hskip 2.84526pt\text{w.p.}\hskip 2.84526pt1-\hat{\mathbb{P}}(a_{t}\succ a^{\prime}_{t})\end{cases}$
The “hat” notation in $\hat{\mathcal{F}}$ and $\hat{\mathbb{P}}$ is to denote corruption. When there is no corruption, we use $\mathbb{P}(a\succ a^{\prime})=\sigma(\mu(a)-\mu(a^{\prime}))$ to denote the probability of event $a\succ a^{\prime}$ and $\mathcal{F}(a,a^{\prime})$ to denote corresponding dueling feedback. It is important to recognize the different roles of the link function $\sigma$ and utility corruption $c_{t}$$\sigma$ captures the inherent (often inevitable) randomness in human choice/behavior even given true utilities (Fishburn et al., 1979; Becker et al., 1963; Bradley and Terry, 1952), whereas our $c_{t}$ is introduced to model inaccurate perception of utilities due to not perfectly knowing her preferences and, importantly, can be alleviated or even eliminated once the user has tried sufficient actions.
Imperfect user feedback generally is not as wild as arbitrary adversarial corruption. A particularly natural pattern of human’s imperfection is that the inaccuracy in our estimations often diminishes as we try more and more actions. That is, humans learn along the way, though how fast the inaccuracy decreases often depends on our learning rate. This hence motivates us to consider the following notion of $\rho$-imperfect user which significantly constrain the corruption sequence $c_{1},c_{2},\cdots$, and force it to diminish at the rate $\rho$.
Definition 1 ($\rho$-Imperfect User).

### v2

@@ Abstract
This paper studies Learning from Imperfect Human Feedback (LIHF), addressing the potential irrationality or imperfect perception when learning from comparative human feedback. Building on evidences that human’s imperfection decays over time (i.e., humans learn to improve), we cast this problem as a concave-utility continuous-action dueling bandit but under a restricted form of corruption: i.e., the corruption scale is decaying over time as $t^{\rho-1}$ for some “imperfection rate” $\rho\in[0,1]$. With $T$ as the total number of iterations, we establish a regret lower bound of $\Omega(\max\{\sqrt{T},T^{\rho}\})$ for LIHF, even when $\rho$ is known. For the same setting, we develop the Robustified Stochastic Mirror Descent for Imperfect Dueling (RoSMID) algorithm, which achieves nearly optimal regret $\tilde{\mathcal{O}}(\max\{\sqrt{T},T^{\rho}\})$. Core to our analysis is a novel framework for analyzing gradient-based algorithms for dueling bandit under corruption, and we demonstrate its general applicability by showing how this framework can be easily applied to obtain corruption-robust guarantees for other popular gradient-based dueling bandit algorithms. Our theoretical results are validated by extensive experiments.
@@ 1 Introduction
Many real-world problems, such as personalized recommendation (Yue and Joachims, 2009; Immorlica et al., 2020; Yao et al., 2022) and fine-tuning generative models (Bai et al., 2022; Casper et al., 2023; Han et al., 2024), require learning from human feedback. Expressing preferences as numerical values or concrete functions is generally challenging for humans. Therefore, an approach that has achieved significant real-world success is to learn humans’ preferences by eliciting their comparative feedback between two options (Bai et al., 2022). A natural theoretical framework capturing such learning task is the seminal dueling bandits framework introduced by Yue and Joachims (2009). The dueling bandit problem features a sequential online learning problem, during which a pair of actions are selected at each round and only their comparison result is revealed to the learner. The comparison outcome between two actions is modeled using a utility function of actions, alongside a link function that determines the probability of each action winning based on their utility difference. This modeling approach is termed as utility-based dueling bandit and achieves great success in generating summaries closely aligned with human preferences (Stiennon et al., 2020).
@@
However, human feedback is often imperfect, a crucial factor that has been largely overlooked in previous studies of dueling bandit. As we are all aware of, humans are not always rational (Posner, 1997) neither perfectly know our preferences (Pu et al., 2012). A powerful framework that can capture this problem is robust bandit learning under adversarial corruption (Bogunovic et al., 2020; Saha and Gaillard, 2022; Di et al., 2024). However, assuming fully adversarial feedback from a regular human seems over-pessimistic at the first glance, hence might make our algorithm development overly conservative. Indeed, ample behavioral studies show that humans often refine their preferences through interactions with the system, navigating a complex tradeoff between exploration and exploitation (Cohen et al., 2007; Wilson et al., 2014). Furthermore, their response errors may exhibit systematic patterns that depend on their past consumption history, rather than merely random noise. Though the sources of imperfections vary a lot in human feedback, one common finding in these studies is that the feedback’s imperfection tends to decrease over time, as humans interact more with the system (Cohen et al., 2007; Wilson et al., 2014; Immorlica et al., 2020; Yao et al., 2022). This premise is the core motivation of this work, which studies dueling bandit learning under corruption that can be arbitrary but have decaying scales.
Our Model and Contributions. We cast Learning from Imperfect Human Feedback (LIHF) as a concave-utility continuous-action dueling bandit learning problem under adversarial utility corruption yet with decaying scale – formally, the scale is upper bounded by $\mathcal{O}(t^{\rho-1})$ at round $t$, for some $\rho\in[0,1]$ (hence the total amount of attack within time $T$ is $\mathcal{O}(T^{\rho})$). A fundamental research question that motivates our study is whether LIHF is fundamentally easier than learning under arbitrarily corrupted feedback.
Our first main technical result hints on a potentially negative answer to the above question. We prove a strong regret lower bound of $\Omega(d\max\{\sqrt{T},T^{\rho}\})$ under LIHF, where $d$ is the action’s dimension. This bound holds even when the decaying rate of imperfection index $\rho$ is known. This lower bound has the same order as recent lower bounds of dueling bandits under (the harder) arbitrary corruption (Agarwal et al., 2021; Di et al., 2024), thus hinting that learning from imperfect human feedback may be no easier than learning from arbitrary adversarial corruption. Notably, however, we use the phrase “hint”, while not an affirmative claim, because our setting with continuous action space and concave utility function is different from the $K$-armed bandits setting of Agarwal et al. (2021) and linear utility function case of Di et al. (2024). Therefore, neither their lower bounds nor their proof techniques are directly applicable to out setting with corruption to strictly concave utilities; see more discussions and comparisons below.
Our second main result is an efficient algorithm with regret upper bound $\tilde{\mathcal{O}}(d\max\{\sqrt{T},T^{\rho}\})$ that matches the above lower bound, up to logrithmic terms, in the same setting (i.e., decaying corruption level $t^{\rho-1}$ with known $\rho$). Unlike previous corruption-robust dueling bandit algorithms based on robust $K$-armed bandits (Agarwal et al., 2021) or robustified LinUCB (Di et al., 2024), our algorithm is a gradient-based algorithm that is built upon the recent Noisy Comparison-based Stochastic Mirror Descent (NC-SMD) of Kumagai (2017). This is natural for our setup with continuous actions and strictly concave utility. While our algorithm itself can be viewed a natural generalization of Kumagai (2017), its analysis is highly non-trivial and much more complex, due to the necessity of accounting for corruption. Indeed, our main technical contribution in this result is the development of a new framework for analyzing the regret of gradient-based algorithms under (possibly arbitrary) corruption. Some novel ingredients of this framework includes a new regret decomposition lemma for dueling bandit which upper bounds total regret by the regret of decision and feedback error due to corruption and a lemma that quantifies bias in gradient estimation. This framework not only yields a tight analysis of our new algorithm, but also easily implies regret bounds for existing algorithms in fully adversarial corruption environments, such as the popular dueling bandit gradient descent (DBGD, Yue and Joachims (2009)). These upper bounds are looser than our lower bound for the LIHF setting. It is an intriguing open question to understand whether these upper bounds under arbitrary corruption are tight, or the lower bound under arbitrary corruption may be stronger than our lower bound under LIHF. Finally, we conduct simulations which validate our theory.
Comparisons with Related Works. Since multiple recent works (Agarwal et al., 2021; Saha and Gaillard, 2022; Di et al., 2024) study dueling bandits under arbitrarily adversarial corruption , it is worthwhile to discuss the key difference between these works and ours. The first key difference is the setting, which leads to fundamentally different algorithms hence analysis. Specifically, our setting with continuous action space and a completely unknown concave utility function is motivated by addressing problems from modern recommender systems where the action (i.e., contents) space is extremely large, not enumerable, thus often embedded as high-dimensional continuous vectors (Chen et al., 2019, 2021). This setting and our techniques are crucially different from earlier works. Agarwal et al. (2021); Saha and Gaillard (2022) consider the $K$-arm dueling bandit setting and utilizes the reduction of Ailon et al. (2014) from dueling bandits to multi-$K$-armed bandits (MAB); (Di et al., 2024) considers linear utility setting where the unknowns is a parameter vector $\theta$, and utilizes techniques from linear contextual bandits based on the optimism principle. The methods all rely on parameter estimation which are not applicable to our setting with arbitrary concave utilities. Moreover, they often need to enumerate all actions to identify the best one under optimism, which are not computationally efficient for our setting with continuous action space. This is why our algorithm is gradient-based and falls within the family of stochastic mirror descent. To the best of our knowledge, this is the first time a corruption-robust algorithm is developed for dueling bandits with continuous actions and strictly concave utilities, which is also why it is necessary for us to develop a new analysis framework. Additionally, we also note that dueling bandits with strictly concave utility and linear utility are generally not comparable, both in terms of their problem difficulties and methodologies. A strictly more general situation is the concave utility case (no need to be strongly concave), however the best upper bound so far for general concave utility is $\mathcal{O}(T^{3/4})$ (Flaxman et al., 2005; Yue and Joachims, 2009), which are considerably worse than the $\mathcal{O}(T^{1/2})$ for linear and for strongly concave utility.
The second main difference between our work and previous is the more restricted class of adversarial environments. Our model is motivated by imperfect human feedback whereas previous studies are motivated by arbitrarily adversarial adversaries. Finally, a more subtle difference is that the corruption in our model is on utility with motivations from learning from imperfect human feedback, whereas corruption in all previous settings directly flip the comparison outcome. These two differences lead to a very different proof techniques of our lower bound. Our proof is more involved because the adversary has limited capability to alter the comparison outcomes. Specifically, the corruption are decaying; moreover, a constant amount of utility corruption only slightly shifts the outcome probability whereas in previous settings each corruption completely changes the outcome.
2 The Problem of Learning from Imperfect Human Feedback
@@
Notation. For a positive integer $T$, we use $[T]$ to denote $\{1,2,\ldots,T\}$. We use standard asymptotic notations including $\mathcal{O}(\cdot)$, $\Omega(\cdot)$, $\Theta(\cdot)$. We use $\tilde{\mathcal{O}}(\cdot)$, $\tilde{\Omega}(\cdot)$, $\tilde{\Theta}(\cdot)$ to hide logarithmic factors. We use $\|\cdot\|_{2}$ to define $L_{2}$ norm, $\|\cdot\|_{\infty}$ to define infinity norm, and $\lambda_{\max}(\cdot)$ to denote maximum eigenvalue.
Motivated by alignment of machine learning models with human preferences (Bai et al., 2022), we study learning from comparative human feedback within the seminal dueling bandit framework (Yue and Joachims, 2009) but account for “imperfections” in human feedback, as described below.
@@
Basics of Continuous Dueling Bandits. We consider the dueling bandit framework with continuous action set $\mathcal{A}\subset\mathbb{R}^{d}$ (Yue and Joachims, 2009). At each round $t\in[T]$, the learner chooses two actions $a_{t},a^{\prime}_{t}$ to present to some human agent, henceforth denoted as the “user”. The user receives utility $\mu(a_{t}),\mu(a^{\prime}_{t})$ from the two actions respectively, and will pick one of these actions following a link function $\sigma(\cdot)$. In the absence of corruption, the user selects action $a_{t}$ with probability $\sigma(\mu(a_{t})-\mu(a^{\prime}_{t}))$ and chooses action $a^{\prime}_{t}$ otherwise. Since the comparison outcome has less errors compared to the exact utility value while still carries useful information about the underlying utility function $\mu$, this form of human feedback has been widely used for learning human preferences (Ailon et al., 2014; Maystre and Grossglauser, 2017; Bengs et al., 2021; Bai et al., 2022). Following standard assumptions in this field (Yue and Joachims, 2009; Kumagai, 2017), we also assume
@@
    the action set $\mathcal{A}$ is a convex compact set that contains the origin, has non-empty interior, and is contained in a $d$-dimensional ball of radius $R$;
@@
    the utility function $\mu:\mathcal{A}\rightarrow\mathbb{R}$ is strongly concave and twice-differentiable. The following constants are useful for our algorithm analysis: $\mu$ is $L^{\mu}$-Lipschitz, $\alpha$-strongly concave,111 $\mu$ is $\alpha$-strongly concave for some $\alpha>0$ if $\mu(x)\geq\mu(y)+\langle\nabla\mu(x),x-y\rangle+\frac{\alpha}{2}\|y-x\|^{2}_{2}$ for any $x,y\in\mathcal{A}$. $\kappa$-smooth, bounded, $R^{\mu}:=\sup_{a,a^{\prime}\in\mathcal{A}}\mu(a)-\mu(a^{\prime})$, and $a^{*}:=\arg\max_{a\in\mathcal{A}}\mu(a)$ is the unique optimal within $\mathcal{A}$;
@@
    the link function $\sigma:\mathbb{R}\rightarrow[0,1]$ is smooth, rotation-symmetric (i.e. $\sigma(x)=1-\sigma(-x)$), and concave for any $x\geq 0$. For the ease of analysis: let $l^{\sigma}_{1}$ [resp. $L^{\sigma}_{1}$] denote the lower [resp. upper] bound of the first-order derivative of $\sigma$. $\sigma$ is $L^{\sigma}$-Lipschitz and its second-order derivative $\sigma^{\prime\prime}$ is $L^{\sigma}_{2}$-Lipschitz and upper bounded by $R^{\sigma}_{2}$.
@@
Some previous works restrict link function to specifically be the logistic link function (Saha, 2021; Xu et al., 2024). This, however, is not needed for our results.
Modeling Imperfect Human Feedback as a Restricted Form of Adversarial Corruption. To incorporate imperfect human feedback, we study a generalization of the dueling bandit framework above, by introducing a corruption term, $c(a,a^{\prime})$, to the utility difference. Formally, let $a\succ a^{\prime}$ represent the event where the user chooses $a$ over $a^{\prime}$. The probability of this event in the presence of corruption $c(a,a^{\prime})$ is denoted by ^$\hat{\mathbb{P}}(a\succ a^{\prime})$, expressed as ^$\hat{\mathbb{P}}(a\succ a^{\prime}):=\sigma(\mu(a)-\mu(a^{\prime})+c(a,a^{\prime}))$. We denote the user’s preferential feedback under corruption as ^$\hat{\mathcal{F}}(a,a^{\prime})$, referred to as imperfect dueling feedback. Mathematically, ^$\hat{\mathcal{F}}(a,a^{\prime})$ follows a binomial distribution with mean ^$\hat{\mathbb{P}}(a\succ a^{\prime})$, as expressed by the following equation.
@@
^$\hat{\mathcal{F}}(a,a^{\prime}):=\begin{cases}1\hskip 2.84526pt\text{w.p.}\hskip 2.84526pt\hat{\mathbb{P}}(a\succ a^{\prime})\\ 0\hskip 2.84526pt\text{w.p.}\hskip 2.84526pt1-\hat{\mathbb{P}}(a\succ a^{\prime})\end{cases}$
The “hat” notation in $\hat{\mathcal{F}}$ and $\hat{\mathbb{P}}$ is to emphasize the presence of corruption. When there is no corruption, we use $\mathbb{P}(a\succ a^{\prime}):=\sigma(\mu(a)-\mu(a^{\prime}))$ to represent the probability of the event $a\succ a^{\prime}$, and $\mathcal{F}(a,a^{\prime})$ to denote the corresponding dueling feedback.
We remark that the term $c(a,a^{\prime})$ above is often called “strong corruption” in the literature of adversarial corruption because it can be introduced after observing actions $a$ and $a^{\prime}$ and the corruption function, $c(\cdot)$, can depend on $\mu$ and past actions (Bogunovic et al., 2020; He et al., 2022; Di et al., 2024). The only difference between our model and the above works on adversarial corruption is a natural restriction to the scale of the corrupted term $c(a,a^{\prime})$. Our motivation is that human feedback, while imperfect, should not be arbitrarily adversarial and tends to improve as human agent interact with the learner. This motivates our following definition of imperfect human feedback.
Definition 1 ($\rho$-Imperfect Human Feedback).